# AgentAda: Skill-Adaptive Data Analytics for Tailored Insight Discovery

### v1

@@ 1 Introduction
A major challenge in advancing LLM-based data analytics is the lack of strong evaluation frameworks that reflect real-world demands. This gap has two key aspects. First, current benchmarks often focus on narrow domains with simple statistical tasks—e.g., Insight-Bench (Sahu et al., 2024) centers on business analytics—overlooking the complexity of broader, real-world analyses. Second, there is no clear way to compare the quality of generated insights. Insight evaluation is subjective and hard to define, and human evaluation, while useful, is difficult to scale due to the expertise required. Progress needs broader, more realistic benchmarks and scalable, expert-informed evaluation methods.
To overcome the first limitation and evaluate the effectiveness of AgentAda we introduce KaggleBench, a benchmark of 700 examples spanning 49 domains and 28 task types. KaggleBench addresses key limitations of prior benchmarks by covering a broader range of analytical tasks that require deeper reasoning and more advanced analytical skills. It provides a more realistic and comprehensive testbed for assessing the capabilities of LLM-based analytics agents. In addition we introduce SCORER (Structured Calibration Of Ratings via Expert Refinement), an LLM-as-a-judge framework guided by human feedback to address the challenge of insight evaluation. While LLM-based evaluation offers scalability, it often lacks grounding in domain-informed quality standards. SCORER bridges this gap by leveraging human-annotated scores, scoring rubrics and sample insights to guide the LLM in evaluating analytical insights in a manner that reflects human judgment. Unlike existing approaches that rely on static prompts or model fine-tuning, SCORER achieves expert-aligned scoring purely through prompt optimization, making it both lightweight and scalable. To our knowledge, this is the first application of prompt-tuned LLM-as-a-judge evaluation for data analytics. In our experiments, we evaluated AgentAda against existing agents (Hu et al., 2024; Sahu et al., 2024; Ge et al., 2023), using KaggleBench as the benchmark and SCORER as the evaluation method. This setup allowed us to assess and improve AgentAda across a wide range of advanced analytical tasks in a reliable and scalable manner.
@@
2 Related Works
We review related work across three areas relevant to our approach: LLM-based data analytics agents, evaluation benchmarks, and LLM-as-a-judge frameworks.
LLM-based Data Analytics. Prior works on LLM-based data analytics agents have explored structured pipelines and multi-agent frameworks, but still face key limitations in adaptability, goal alignment, efficiency, and generalization. Multi-agent systems (Rasheed et al., 2024; Fischer & Biemann, 2024; Chugh et al., 2023) break down problems into sub-tasks handled by specialized agents. But they lack guidance in choosing the right analytical methods, often resulting in basic and shallow insights. They also struggle to adapt to specific user goals or personas. Other systems like InfiAgent (Hu et al., 2024) and Data Interpreter (Hong et al., 2024) use strategies like ReAct (Yao et al., 2023) and hierarchical modeling to generate structured code, but without grounding in task context or access to skill-specific usage examples, their outputs are often error prone and rely heavily on inefficient iterative debugging. In contrast, AgentAda’s skill-informed pipeline enables efficient, goal-driven, and adaptable analysis, which generalizes across various tasks and domains.
@@ 2 Related Works
3
KaggleBench – A Data Analytics Benchmark
KaggleBench is a curated benchmark designed to evaluate the analytical capabilities of data analytics agents across a wide range of tasks, skills, and domains.
Below, we outline the data collection and construction process in detail. See Appendix A for statistics on KaggleBench.
3.1 Dataset Notebooks QA Generation
The dataset is sourced from high-quality Jupyter notebooks published by by data analysts on Kaggle222https://www.kaggle.com, a popular platform for data science and analysis. We collected 700 notebooks covering diverse analytical domains and tasks types. Each notebook contains structured workflows, markdown summaries, and corresponding datasets, making them well-suited for insight-focused evaluation. To create fine-grained QA examples, we parsed notebooks into cell batches and used GPT-4o to: (I) generate QA pairs, and (II) assign each question a task and skill label from a predefined library. The full list of supported tasks and analytical skills is provided in Appendix B. Answers were extracted from human-written markdown conclusions or code cell outputs. This ensured that the QA pairs reflected the actual reasoning and results within the notebook. We then evaluated the source of the answers (i.e., whether they originated from markdown or code cell outputs) using a simple RAG-Token Model introduced by Lewis et al. (2020). QA pairs with invalid tasks or skills were subsequently filtered out (12.28% percent of questions were removed). We then used an LLM to select the top well-framed QA pairs from each notebook based on diversity. The prompts used are available OpenAI et al. (2024) in Appendix E.1.
3.2 Goal and Persona Generation
To support goal- and persona-aware evaluation of analytical insights, we generated a corresponding goal and persona for each notebook in KaggleBench. The goal is a concise statement capturing the purpose of the analysis of the notebook, focusing on what and why something is being analyzed, without specifying how the analysis is performed. The persona describes the role or perspective (e.g., data analyst, business strategist) from which the analysis is conducted. An example of a generated goal and persona is shown in Figure 1. Note that while a dataset may support multiple analytical directions, we extract the goal and persona that reflect the specific analysis actually carried out in the notebook. Both were extracted using GPT-4o, with the prompting strategy detailed in Appendix E.1.
@@ 4 AgentAda – A Skill-Informed Data Analytics Agent
4.1 Skill Set Collection
We curated a library of diverse data analytics skills, covering a range of tasks and algorithms as listed in Table 7. These skills, implemented in Python, were primarily sourced from Kaggle notebooks. Then, they are organized using a standardized workflow that includes data preparation, modeling, evaluation, and visualization. For each skill, we also extracted a concise summary to capture its core functionality and intended use.
4.2 Dual Stage Advanced Question Generation
@@
Insight generation begins with asking the right questions. To guide AgentAda in producing meaningful, goal-aligned insights for a given dataset, we start with designing a two-stage question generation process. In the first stage, we generate a set of basic data analytics questions using the dataset, goal, and persona. We focus on straightforward tasks such as filtering or simple aggregations. In the second stage, we use the available skills in the skill library along with the generated simple questions to generate more advanced questions that require complex reasoning and advanced techniques to analyze. This setup helps AgentAda uncover deeper patterns in the data. Some examples of both basic and advanced questions, along with corresponding analyses, are provided in Appendix E.2. Detailed prompts for both stages are provided in Appendix E.2.
4.3 Category prediction
@@
To evaluate the performance of AgentAda against other analytics agents, it is important that the insights being compared are organized around similar high-level themes. To support this, we introduce an insight category prediction module that estimates the overarching analytical themes likely to emerge from the responses to each set of questions. We achieve this by prompting GPT-4o with the dataset description, analysis goal, and the list of generated questions, asking it to predict the top three insight categories that will likely capture the essence of the responses. More details about the prompting strategy for this module are provided in Appendix E.3.
4.4 Skill Matcher
@@
For each question in the advanced set, we retrieve the most relevant analytical skill from the skill library using a Hybrid Retrieval-Augmented Generation (RAG) system (Dong et al., 2024; Li et al., 2024; Su et al., 2024; Shi et al., 2024; Sticha, 2023). This system connects natural language questions to executable analysis by combining semantic search with structured mappings between skill descriptions and their corresponding implementations. The skill matcher helps guide AgentAda’s analysis toward the most suitable techniques for answering each question accurately and efficiently. It operates in four steps: (I) an LLM interprets the question to identify its analytical intent and underlying task type, (II) the question is embedded and matched against skill summaries using OpenAI embeddings (Neelakantan et al., 2022), (III) the top- most relevant skills are retrieved from the library ( in our setup) and (IV) the selected skill, including its summary and implementation, is passed to the code generation module to guide the next stage of analysis. The full prompting strategy for the skill matcher is provided in Appendix E.4.
4.5 Code Generation
@@
After retrieving the question and relevant skill, the code generation module produces structured, executable code to meet the analytical goal. It takes the data schema, question, predicted skill, and its summary as input to generate code with visualizations and key statistics. The skill guides the LLM in producing clear, complete code for preprocessing, analysis, plotting, and metric extraction. If execution fails, the error message is added to the prompt for regeneration—up to three attempts per question—enabling self-correction without manual input. On average, we observe generations per dataset with 5 questions. Full prompting details are available in Appendix E.5.
4.6 Answer Generation
@@
The next step is to generate responses from the plots and statistical outputs generated by the executed code for each question generated by the question generation module 4.2. For this we use a multimodal LLM, that takes as input the question, generated plot, and key statistics to produce a structured response. These responses are then summarized into concise bullet points for clarity and ease of interpretation. Both the answer generation prompts are provided in Appendix E.6.
4.7 Insight Generation
@@
In the final insight extraction step, we aggregate the answers across all questions for a dataset and prompt a strong LLM with the dataset description, overall goal, generated answers, and insight categories to produce goal-aligned and actionable insights. Prompting strategies for this stage is detailed in Appendix E.7.
@@ 5 SCORER
We evaluate the quality of generated insights using SCORER (Structured Calibration Of Ratings via Expert Refinement), a LLM-as-a-judge framework that aligns the scores of the evaluator model with human judgment through prompt optimization rather than model fine-tuning. Our central hypothesis is that an LLM can approximate expert evaluation if provided with the right contextual cues derived from human ratings and rationales. Instead of manually encoding examples or explanations into the prompt, SCORER automatically learns how humans score by optimizing the prompt itself. To create SCORER, first, we define a shared set of evaluation criteria as in Section 6.1 used consistently by both human evaluators and SCORER. Then, we formulate context extraction as a prompt optimization problem. The optimizer (Yuksekgonul et al., 2024) is initialized with a starter prompt that enables the LLM to score insights independently. Then the optimizer iteratively refines the prompt by minimizing the mean squared error (MSE) between LLM’s scores and human scores. The final expert aligned prompt closely mirrors expert evaluation patterns and helps the model to score insights in a human-aligned manner. This approach retains the scalability of LLM-based evaluation, while significantly improving alignment with expert judgment. Full details of the starter prompt and expert aligned prompt are provided in Appendix F. As shown in Figure 3, SCORER’s the validation loss decreases over steps, indicating improved alignment with human evaluation results.
@@ 6 Experiments & Results
All LLM interactions, including skill retrieval, code generation, and insight extraction, were performed via API calls to OpenAI’s GPT-4o (OpenAI et al., 2024) and text-embedding-3-small models.
@@ 6.1 Evaluation of AgentAda’s Skill Abilities
We evaluate the quality of insights generated by AgentAda by comparing it against several existing analytics agents, including Poirot (Sahu et al., 2024), Pandas AI (Fischer & Biemann, 2024), InfiAgent (et al., 2024), MetaGPT Ge et al. (2023), and direct prompting with GPT-4o. Also, to validate our core hypothesis that skill-informed agents produce deeper insights, we also introduce a variant of AgentAda that removes skill guidance. In this baseline, the LLM attempts to infer and apply the appropriate skill without access to the curated skill library. We denote this setup as W/O Skill, while the full version of AgentAda is referred to as W Skill in our results. We compare the performance of AgentAda’s W Skill and W/O Skill variants across six rubrics: depth of analysis, relevance to goal, persona consistency, coherence, answering question adequately, and plot conclusion quality. Instead of evaluating the final dataset-level insights, we assess the quality of individual answers produced for each question in the answer generation stage 4.6. This is because both variants use the same prompt and LLM for final insight generation, which often leads to similar-looking outputs. Since that can mask the real differences between the two setups and confuse human evaluators, we perform the comparison directly on the question-answer pairs to better capture the effect of skill guidance. We compared the performance of AgentAda against other agents across all rubrics except answering question adequately, as where we conducted the evaluation at the final insight level instead of individual answers and there were no same questions to evaluate in this criterion.
@@
We conducted a human evaluation of the W Skill and W/O Skill variants of AgentAda on 100 datasets spanning diverse analytical tasks and domains. The evaluation was split into 10 batches of 10 questions, each reviewed by three independent annotators with data analytics backgrounds to ensure consistency and assess analytical depth, relevance, and reasoning quality. We used Fleiss’ Kappa (Fleiss, 1971) to measure annotator agreement and assess the reliability of our evaluation (Table 1). Most criteria showed strong agreement, while Goal Relevance and Persona Consistency had lower scores—expected given their subjective nature. Annotators may avoid the “Tie” option in borderline cases, adding noise, and assessing persona alignment often depends on individual interpretation of tone and perspective. The human evaluation setup is described in detail in Appendix C. Additional statistics on the human evaluators are provided in D.
| 0.8842 | 0.8297 | 0.8431 | 0.7658 | 0.8274 | 0.8765 |
@@
Table 2 summarizes the results of our human evaluation. The scores provided mean amongst all the submitted feedbacks, in how many cases of 4 possible options selected (win with skill, win without skill, tie, and none). Based on the results, the skill-informed version of AgentAda outperforms the W/O skill across all rubrics. Notably, the biggest margin is observed in Depth of Analysis, confirming our hypothesis that retrieved skills lead to deeper insights. Rubrics like Relevance to Goal and Persona Consistency show a high number of ties (49.22% and 61.44%), reflecting more subtle differences between the two variants, which aligns with their lower Fleiss’ kappa scores, indicating greater subjectivity. Refer to Appendix G for qualitative analysis.
@@
We evaluated AgentAda using SCORER on KaggleBench containing 700 datasets spanning diverse analytical tasks. To train SCORER, we first collected human evaluation scores on datasets and split them into a train-test split. Then, the starter prompt was optimized using TextGrad (Yuksekgonul et al., 2024) to minimize the mean squared error (MSE) between the LLM-predicted scores and human evaluation scores. After optimization, the SCORER prompt achieved a validation loss of , indicating strong alignment with human judgment and reliable replication of expert preferences. We used the optimized human aligned prompt to score insights across all datasets in KaggleBenchand compare AgentAda against other baselines. The results comparing the skill-informed (overall and top-5 frequent tasks) variant of AgentAda and without skill variant is presented in Table 3. The most significant gains are observed in Depth of Analysis and Coherence, where over of the responses are rated better when guided by retrieved skills. This supports our core hypothesis. On execution-aligned rubrics like Answers Question Adequately and Plot Conclusion, the skill-informed model again performs better. This shows that guided code generation helps generate complete responses and stronger visual reasoning. Also, it is worth noting that these findings closely mirror trends observed in our human evaluation (Table 2), with high alignment across most rubrics. This further validates SCORER’s effectiveness as a lightweight and scalable proxy for human judgment.
@@
We also compared AgentAda with other baseline agents, the results of which are presented in Table 4. Across all criteria, AgentAda consistently outperforms all baselines, confirming the effectiveness of skill-informed analysis. Notably, AgentAda shows the strongest performance gains over the Pandas agent, with win rates of in Depth of Analysis, in Coherence, and in Plot Conclusion, indicating a clear advantage in generating deeper, clearer, and more structured insights. This performance gap stems from the design of the Pandas Agent, which relies on rule-based natural language to code translation while AgentAda supports skill-guided code generation which result in richer and deeper responses. AgentAda also demonstrates strong performance against powerful agents like GPT-4o and MetaGPT. Though these models are capable of generic reasoning, their lack of analytical skill grounding leads to shallow insights. Overall, these findings reinforce the value of incorporating structured analytical skills into LLM-based data agents.

### v2

@@ 1 Introduction
A major challenge in advancing LLM-based data analytics is the lack of strong evaluation frameworks that capture real-world demands. Two gaps stand out. First, current benchmarks often focus on narrow domains with simple statistical tasks (e.g., Insight-Bench (Sahu et al., 2024) focuses on business analytics) and fail to reflect the complexity of broader analytical settings. Second, there is no clear way to compare the quality of generated insights. Insight evaluation is subjective and hard to define. Human evaluation, while useful, is difficult to scale due to the expertise required. Progress needs broader, more realistic benchmarks and scalable, expert-informed evaluation methods.
To overcome the first limitation and evaluate the effectiveness of AgentAda we introduce KaggleBench, a benchmark of 700 examples spanning 49 domains and 28 task types. KaggleBench addresses key limitations of prior benchmarks (Sahu et al., 2024) by covering a broader range of analytical tasks that require deeper reasoning and more advanced analytical skills. It provides a more realistic and comprehensive testbed for assessing the capabilities of LLM-based analytics agents. In addition we introduce SCORER (Structured Calibration Of Ratings via Expert Refinement), an LLM-as-a-judge framework guided by human feedback for evaluating analytical insights. Unlike prior approaches that rely on static prompts or fine-tuning, SCORER uses prompt optimization with human-annotated rubrics to achieve expert-aligned scoring, making it lightweight and scalable. To our knowledge, this is the first application of prompt-tuned LLM-as-a-judge evaluation in data analytics. In our experiments, we benchmarked AgentAda against existing agents (Hu et al., 2024; Sahu et al., 2024; Ge et al., 2023) on KaggleBench, with SCORER as the evaluation method.
@@
2 Related Work
LLM-based Data Analytics. Prior works on LLM-based data analytics agents have explored structured pipelines and multi-agent frameworks, but still face key limitations in adaptability, goal alignment, efficiency, and generalization. Multi-agent systems (Rasheed et al., 2024; Fischer & Biemann, 2024; Chugh et al., 2023) break down problems into sub-tasks handled by specialized agents. But they lack guidance in choosing the right analytical methods, often producing surface-level summaries or descriptive statistics rather than deeper diagnostic or prescriptive insights. They also struggle to adapt to specific user goals or personas. Other systems like InfiAgent (Hu et al., 2024) and Data Interpreter (Hong et al., 2024) use strategies like ReAct (Yao et al., 2023) and hierarchical modeling to generate structured code. But without incorporating the specific analytical objectives, dataset characteristics, or examples of how domain-relevant skills should be applied, their outputs are often error prone and rely heavily on inefficient iterative debugging. In contrast, AgentAda’s skill-informed pipeline enables efficient, goal-driven, and adaptable analysis, which generalizes across various tasks and domains.
@@ 2 Related Work
3 KaggleBench –Data Analytics Benchmark
KaggleBench is a curated benchmark designed to evaluate the analytical capabilities of data analytics agents across a wide range of tasks, skills, and domains. Below, we outline the data collection and construction process in detail. See Appendix A for statistics on KaggleBench.
Dataset Notebooks QA Generation. The dataset is sourced from high-quality Jupyter notebooks published by data analysts on Kaggle222https://www.kaggle.com , a leading platform for data science and analysis. We collected 700 notebooks spanning diverse analytical domains and task types. Each notebook contains structured workflows, markdown summaries, and datasets, making them well-suited for insight-focused evaluation. To construct fine-grained QA examples, we parsed notebooks into cell batches and used GPT-4o to (I) generate QA pairs and (II) assign each question a task and skill label from a predefined library (Appendix B). Answers were drawn directly from markdown conclusions or code outputs, ensuring that QA pairs reflect the actual reasoning and results in the notebooks. We further verified the answer source (markdown vs. code) using a RAG-Token Model (Lewis et al., 2020). This RAG-based grounding helps maintain factuality, as QA pairs are generated directly from notebook content, making them reliable for evaluating whether models produce factually correct insights. QA pairs with invalid tasks or skills were filtered out (12.28% removed), and an LLM was used to select the top diverse, well-framed QA pairs from each notebook. The prompts are provided in Appendix I.1 (OpenAI et al., 2024).
Goal and Persona Generation. To support goal- and persona-aware evaluation of analytical insights, we generated a corresponding goal and persona for each notebook in KaggleBench. The goal is a concise statement capturing the purpose of the analysis of the notebook, focusing on what and why something is being analyzed, without specifying how the analysis is performed. The persona describes the role or perspective (e.g., data analyst, business strategist) from which the analysis is conducted. An example of a generated goal and persona is shown in Figure 1. Note that while a dataset may support multiple analytical directions, we extract the goal and persona that reflect the specific analysis actually carried out in the notebook. Both were extracted using GPT-4o, with the prompting strategy detailed in Appendix I.1.
@@ 4 AgentAda – A Skill-Informed Data Analytics Agent
Skill Set Collection. We curated a library of diverse data analytics skills, covering a broad range of tasks and algorithms 7. Implemented in Python, these skills were primarily sourced from Kaggle notebooks. To build the library, we identified high-quality notebooks across domains, converted them to markdown. Here, we retained only workflow-relevant code blocks for data preparation, modeling, evaluation, and visualization, discarding exploratory or environment-specific cells. Next, these curated code workflows were provided as input to an LLM to generate concise, text-based descriptions for each workflow. Importantly, these descriptions were not intended to merely summarize the code. They serve as the skill itself, a knowledge base that clearly explains what the algorithm does, when it should be applied, and how it can be used to solve a data analysis problem. Also, because the library is organized around modular, reusable workflows, it is naturally extensible, and new skills can be easily added as data analysis practices evolve.
Although both the skill library and KaggleBench are sourced from Kaggle, there is no data leakage between them. The skill library captures reusable algorithmic workflows and their descriptions, while KaggleBench independently constructs QA pairs from datasets. This separation ensures that the skills represent general methods, whereas the benchmark evaluates their application on unseen tasks and data.
@@
Dual Stage Advanced Question Generation. Insight generation begins with asking the right questions. To guide AgentAda in producing meaningful, goal-aligned insights for a given dataset, we start with designing a two-stage question generation process. In the first stage, we generate a set of basic data analytics questions using the dataset, goal, and persona. We focus on straightforward tasks such as filtering or simple aggregations. In the second stage, we use the available skills in the skill library along with the generated simple questions to generate more advanced questions that require complex reasoning and advanced techniques to analyze. This setup helps AgentAda uncover deeper patterns in the data. Some examples of both basic and advanced questions, along with corresponding analyses, are provided in Appendix I.2. Detailed prompts for both stages are provided in Appendix I.2.
@@
Category prediction. To evaluate the performance of AgentAda against other analytics agents, it is important that the insights being compared are organized around similar high-level themes. To support this, we introduce an insight category prediction module that estimates the overarching analytical themes likely to emerge from the responses to each set of questions. We achieve this by prompting GPT-4o with the dataset description, analysis goal, and the list of generated questions, asking it to predict the top three insight categories that will likely capture the essence of the responses. More details about the prompting strategy for this module are provided in Appendix I.3.
@@
Skill Matcher. For each question in the advanced set, we retrieve the most relevant analytical skill from the skill library using a Hybrid Retrieval-Augmented Generation (RAG) system (Dong et al., 2024; Li et al., 2024; Su et al., 2024a; Shi et al., 2024; Sticha, 2023). This system connects natural language questions to executable analysis by combining semantic search with structured mappings between skill descriptions and their corresponding implementations. The skill matcher helps guide AgentAda’s analysis toward the most suitable techniques for answering each question accurately and efficiently. It operates in four steps: (I) an LLM interprets the question to identify its analytical intent and underlying task type, (II) the question is embedded and matched against skill summaries using OpenAI embeddings (Neelakantan et al., 2022), (III) the top- most relevant skills are retrieved from the library ( in our setup) and (IV) the selected skill, including its summary and implementation, is passed to the code generation module to guide the next stage of analysis. The full prompting strategy for the skill matcher is provided in Appendix I.4.
@@
Code Generation. After retrieving the question and relevant skill, the code generation module produces structured, executable code to meet the analytical goal. It takes the data schema, question, predicted skill, and its summary as input to generate code with visualizations and key statistics. The skill guides the LLM in producing clear, complete code for preprocessing, analysis, plotting, and metric extraction. If execution fails, the error message is added to the prompt for regeneration, up to three attempts per question. This enables self-correction without manual input. On average, we observe generations per dataset with 5 questions. Full prompting details are available in Appendix I.5.
@@
Answer Generation. The next step is to generate responses from the plots and statistical outputs generated by the executed code for each question generated by the question generation module. For this we use a multimodal LLM ((OpenAI et al., 2024)), that takes as input the question, generated plot, and key statistics to produce a structured response. These responses are then summarized into concise bullet points for clarity and ease of interpretation. Both answer generation prompts are provided in Appendix I.6.
@@
Insight Generation. In the final insight extraction step, we aggregate the answers across all questions for a dataset and prompt the LLM with the dataset description, overall goal, generated answers, and insight categories to produce goal-aligned and actionable insights. Prompting strategies are detailed in I.7.
@@ 5 SCORER
We evaluate the quality of generated insights using SCORER (Structured Calibration Of Ratings via Expert Refinement), an LLM-as-a-judge framework that aligns model scores with human judgment through prompt optimization rather than fine-tuning. The key idea is that an LLM can approximate expert evaluation when guided by contextual cues from human ratings and rationales. SCORER helps evaluate on a shared set of evaluation criteria (Section 6.1) and formulates prompt refinement as an optimization problem using TextGrad (Yuksekgonul et al., 2024). It minimizes the mean squared error between LLM and human scores. Starting from a simple starter prompt, the optimizer iteratively refines toward an expert-aligned prompt that mirrors human scoring patterns.
This approach preserves the scalability of LLM-based evaluation and improves reliability and alignment with human judgment. Full prompt details are provided in Appendix J. Figure 3 shows validation loss decreasing over optimization steps, indicating closer agreement with human evaluators. More results on robustness and generalizability of SCORER is in Appendix L.
@@ 6 Experiments & Results
All LLM interactions were performed via API calls to OpenAI’s GPT-4o (OpenAI et al., 2024) and text-embedding-3-small models.
@@ 6.1 Evaluation of AgentAda’s Skill Abilities
We evaluate the quality of insights generated by AgentAda against several analytics agents, including Poirot (Sahu et al., 2024), Pandas AI (Fischer & Biemann, 2024), InfiAgent (Hu et al., 2024), MetaGPT (Ge et al., 2023), and direct prompting with GPT-4o. To test our core hypothesis that skill-informed agents produce deeper insights, we also include a variant of AgentAda without skill guidance. In this baseline (W/O Skill), the LLM infers and applies skills without access to the curated skill library, while the full version is denoted W Skill. We compare all these agents across six rubrics: depth of analysis, relevance to goal, persona consistency, coherence, answering question adequately, and plot conclusion quality (definitions in Appendix C). When comparing the W/O Skill and W Skill variants, instead of evaluating only final dataset-level insights, we assess the quality of individual answers in the answer generation stage. This avoids the masking effect of similar-looking final outputs, which use the same LLM and prompt in both variants, and allows us to more directly capture the impact of skill guidance. Performance against other agents is reported across all rubrics, with the exception of answering question adequately, which was evaluated at the final insight level where questions did not align across systems.
@@
In the first experiment we conducted a human evaluation of the W Skill and W/O Skill variants of AgentAda on 100 datasets spanning diverse analytical tasks and domains. The evaluation was split into 10 batches of 10 questions, each reviewed by three independent annotators with data analytics backgrounds to ensure consistency and assess analytical depth, relevance, and reasoning quality. We used Fleiss’ Kappa (Fleiss, 1971) to measure annotator agreement and assess the reliability of our evaluation (Table 8). Most criteria showed strong agreement, while Goal Relevance and Persona Consistency had lower scores, expected given their subjective nature. Annotators may avoid the “Tie” option in borderline cases, adding noise, and assessing persona alignment often depends on individual interpretation of tone and perspective. The human evaluation setup is described in detail in Appendix C. Additional statistics on the human evaluators are provided in D.
@@
Table 2 summarizes the results of our human evaluation. Scores represent the distribution of judgments across four options: win with skill, win without skill, tie, and none. Overall, the skill-informed version of AgentAda outperforms the W/O Skill variant across all rubrics, with the largest margin in Depth of Analysis, supporting our hypothesis that retrieved skills enable deeper insights. In contrast, rubrics such as Relevance to Goal and Persona Consistency show a high proportion of ties (49.22% and 61.44%), highlighting more subtle differences between the variants. This is consistent with their lower Fleiss’ kappa scores, which indicate greater subjectivity.
To better understand the cases where the W/O Skill variant was judged superior, we conducted a qualitative study (Table 1). We find that when skilled responses are preferred, they typically apply the correct analytical method (e.g., PCA, collaborative filtering, market basket analysis), while unskilled answers rely on shallow heuristics such as simple counting. In the cases where unskilled answers were preferred, the skilled variant introduced unnecessary complexity. (e.g., using PageRank instead of counting connected components). Importantly, these skilled responses were not incorrect. They often applied valid analytical methods, but their complexity was not required for the question at hand, making the simpler unskilled answers more appealing. Additional qualitative analysis is provided in Appendix M.
| Task | Question | Preferred | Unskilled Answer | Skilled Answer |
Recommendation
Systems |
What trends are observed in user preferences across different categories, and how can this guide targeted recommendations? | Unskilled | Electronics and Home Appliances are the most frequently chosen categories. These can be targeted with personalized promotions. | Using Collaborative Filtering, we find latent factors influencing user-category affinity scores. The model predicts cross-domain co-preference between Electronics and Smart Gadgets. |
| Market Analysis | What purchase patterns were observed based on different days of the week? | Unskilled | Weekends show higher sales of snacks and beverages, while weekdays focus on essentials like milk and bread. | Association rule mining identifies high-confidence weekend-specific rules such as {chips} {soda}, but this complicates a simple weekday/weekend pattern. |
Recommendation
Systems |
How do the preferences of users who rate niche items overlap with general user preferences, and how can this guide recommendations? | Skilled | Users who rate niche items also seem to buy from common categories like Electronics. | Collaborative Filtering reveals that niche users share latent factors with mainstream users, allowing cross-category recommendations such as niche Book readers also preferring certain Gadgets. |
| Topic Modeling | What are the topics identified in the articles by matrix factorization? | Skilled | From word counts, I can guess the main categories are business, politics, and sports. | Latent Semantic Analysis identifies five topics with top terms: (1) business/finance, (2) tech/innovation, (3) politics/government, (4) sport/teams, (5) entertainment/media. |
Anomaly
Detection |
How did adjusting the contamination parameter affect anomaly detection results? | Skilled | When we increased the parameter, we saw more anomalies flagged, but it’s hard to quantify. | Increasing contamination from 0.02 to 0.06 raised detected anomalies by 48% but also increased false positives, showing a precision–recall tradeoff. |
| Market Analysis | Which product combination shows the strongest lift score? | Skilled | Customers often buy pasta with pasta sauce, so they likely have a strong relationship. | Market Basket Analysis shows Pasta + Pasta Sauce has the strongest lift of 2.8, meaning they are nearly 3x more likely to be bought together than by chance. |
@@
We evaluated AgentAda using SCORER on KaggleBench containing 700 datasets spanning diverse analytical tasks. To train SCORER, we first collected human evaluation scores on datasets and split them into a train-test split. Then, the starter prompt was optimized using TextGrad (Yuksekgonul et al., 2024) to minimize the mean squared error (MSE) between the LLM-predicted scores and human evaluation scores. After optimization, the SCORER prompt achieved a validation loss of , indicating strong alignment with human judgment and reliable replication of expert preferences. More details on SCORER human alignment is provided in Appendix H. We used the optimized human aligned prompt to score insights across all datasets in KaggleBenchand compare AgentAda against other baselines. The results comparing the skill-informed (overall and top-5 frequent tasks) variant of AgentAda and without skill variant is presented in Table 3. The most significant gains are observed in Depth of Analysis and Coherence, where over of the responses are rated better when guided by retrieved skills. This supports our core hypothesis. On execution-aligned rubrics like Answers Question Adequately and Plot Conclusion, the skill-informed model again performs better. This shows that guided code generation helps generate complete responses and stronger visual reasoning. Also, it is worth noting that these findings closely mirror trends observed in our human evaluation (Table 2), with high alignment across most rubrics. This further validates SCORER’s effectiveness as a lightweight and scalable proxy for human judgment.
@@
We compared AgentAda with baseline agents, as shown in Table 4. Across all criteria, AgentAda consistently outperforms all baselines, confirming the effectiveness of skill-informed analysis. Notably, AgentAda shows the strongest performance gains over the Pandas agent, with win rates of in Depth of Analysis, in Coherence, and in Plot Conclusion, indicating a clear advantage in generating deeper, clearer, and more structured insights. This gap reflects the limitations of Pandas, which relies on rule-based natural language–to–code translation, compared to AgentAda’s skill-guided code generation. AgentAda also demonstrates strong performance against powerful agents like GPT-4o and MetaGPT. Though these models are capable of generic reasoning, their lack of analytical skill grounding leads to shallow insights. We also benchmarked AgentAda against code generation agents like ECA and Evor (Wang et al., 2024; Su et al., 2024b) with details provided in Appendix E and assessed the performance in other benchmarks in Appendix K. Overall, these findings reinforce the value of embedding structured analytical skills into LLM-based data agents.
6.3 Evaluating the Factuality of the Insights
We assess the factual consistency of agent outputs using multiple factuality metrics. Proprietary LLMs such as GPT-5 OpenAI (2025), GPT-4o OpenAI et al. (2024), and LLaMA-3.3-70B Grattafiori et al. (2024) were prompted to rate outputs on a – scale, where 1 denotes frequent errors and 5 denotes full consistency with the data. In addition, we apply FactScore Min et al. (2023), which decomposes outputs into atomic claims and verifies them against ground-truth tables in KaggleBench. In our setup, we follow the original FactScore methodology and use the KaggleBench dataset, consisting of 700 examples across diverse notebooks. For each notebook, we utilize the associated QA list to extract evaluation questions and restrict atomic claims to those grounded in the underlying data tables, excluding suggestions or speculative outputs not present in the source. We compute FactScore with multiple verifiers, including GPT-5, GPT-4o, LLaMA-3.3-70B, and Claude-Opus 4. Table 11 shows that AgentAda with skills consistently outperforms its ablation and other baselines, confirming that skill retrieval improves factual grounding.
6.4 Evaluating the correctness of insights
Next, we evaluate the correctness of analysis by AgentADA and compare it with different baselines on KaggleBench. We report accuracy against ground-truth QA and the other rubrics (Depth of Analysis, Answers Adequately, Coherence, Relevance to Goal, Persona Consistency) are rated on a 1–5 scale using SCORER. As shown in Table 5, AgentAda with skill retrieval consistently achieves the highest scores, outperforming both its w/o-skill variant and all other agents. The results show that selecting more appropriate skills leads to stronger analysis and, in turn, more reliable insights.
| Rubric | AgentAda w skill | w/o skill | Poirot | Pandas | InfiAgent | Meta GPT | GPT-4o |
| Accuracy | 90.6 | 82.6 | 81.2 | 80.4 | 84.2 | 85.5 | 78.8 |
| Depth of Analysis | 4.46 | 4.10 | 4.00 | 3.97 | 4.06 | 4.03 | 3.92 |
| Answers Adequately | 4.41 | 4.06 | 3.95 | 3.91 | 4.02 | 3.99 | 3.88 |
| Coherence | 4.48 | 4.09 | 3.97 | 3.94 | 4.03 | 4.00 | 3.90 |
| Relevance to Goal | 4.44 | 4.11 | 3.98 | 3.95 | 4.05 | 4.02 | 3.89 |
| Persona Consistency | 4.42 | 4.07 | 3.96 | 3.93 | 4.04 | 4.01 | 3.87 |